A Bibliography of Publications in Theoretical Computer Science: 2025–2029

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

15 January 2025 Version 1.03

Title word cross-reference

0, 1, 2 [45]. b [36]. d [36]. k [45, 62, 77].

-ary [45]. -center [77]. -error [45]. -way [62].

1 [21, 54]. 12 [57, 43]. 19 [70].

2 [56]. 2025 [56, 57, 70, 21, 43]. 2d [64].

943 [39].

ABCD [68]. accident [8]. action [53]. acyclic [9]. Adaptive [66, 23]. Adding [54]. agent [72, 7]. agreement [48]. Algebraic [38]. Algorithm [1, 39, 8]. Algorithms [9, 65, 25, 32, 62, 77, 30, 37, 75, 10]. allocations [67, 69]. analog [53]. analysis [17]. anonymous [52, 23]. any [48]. approach [69]. Approximating [50, 6, 2]. Approximation [25, 10, 60, 65, 79]. area [76]. area-proportional [76]. arithmetic [53]. arrangement [34]. ary [45]. assisted [5]. asynchronous [44]. asynchrony [32]. attribute [5]. attribute-based [5]. Availability [1, 39]. axis [48].

Balancing [59]. base [36]. Based [1, 39, 66, 40, 69, 5]. Big [37]. bijective [73]. black [16, 78]. blocking [8]. Board [20, 42, 55, 58, 71]. bounded [77]. box [78]. bubble [17]. bubble-sort [17]. Burrows [73].

candidate [31]. capacitated [79]. cardinality [49]. center [77]. central [51]. chains [3]. charging [16]. Chemical [53]. chordal [4]. chores [69]. ciphertext [74]. circle [52, 10]. classes [13]. classification [37]. Cloud [1, 39, 5]. cloud-assisted [5]. clustering [79, 77]. collective [23]. coloring [4]. communities [68]. Comparison [37]. complete [47]. Complexity [63, 67]. compressed [74]. computations [53]. computer [39]. computers [53]. computing [44]. concerning [24, 63]. conditional [34]. Conjecture [54]. connectivity [34]. constraint [49]. constraints [54]. Constructing [16]. contact [76]. containing [10]. containment [61]. contextual [24]. contiguous [65]. counting [29]. cover [2]. covering [25]. cubes [47]. cut [27].

data [59, 30, 37]. decision [24, 6]. declined [17]. delays [62]. deletions [24]. delivery [7]. Density [45]. Deterministic [62]. Dichotomies [61]. dichotomy [15]. differential [30]. directed [9]. direction [54]. disjunctive [22]. disks [25]. distance [65]. Distributed [44, 66, 77]. diversity [12, 59]. division [36]. domain [15]. domination [63]. Double [17, 27]. double-cut-and-joins [27]. DQN [75]. dual [62]. during [8]. Dynamic [7, 23]. Dynamically [22]. dynamics [72].

e-healthcare [5]. Editorial [20, 42, 55, 58, 71]. Efficient [8]. EFX [69]. Egalitarian [67]. Enabling [34]. encrypt [74]. enhancement [74]. equal [73]. equal-letter [73]. equations [30]. equivalent [22]. error [45]. Euclidean [36]. evolutionary [72]. Exact [65, 64, 35]. exchange [30]. existence [13]. Extremal [3].

facility [31]. failure [48]. fast [78]. Fault [48, 17]. Fault-tolerant [48]. February [56, 57, 70]. finder [78]. Finding [29, 7, 10]. fine [60]. fine-grained [60]. Finer [60].

Finer-grained [60]. formalization [30]. FPT [79]. freshness [59]. funneled [13].

Gathering [52, 14]. genomes [27].

Geography [19]. gradient [72]. grained [60]. graph [34]. graphs [9, 63, 4]. grid [14]. grids [18].

hardness [60]. healthcare [5]. hiding [5]. high [34]. Holant [15]. hypotheses [6].

Implementing [53]. Improved [75]. independent [47]. independently [28]. index [3]. indivisible [69]. inexact [64]. infinite [18, 14]. information [40, 5]. information-based [40]. Injective [4]. insertions [24]. Instance [40]. Instance-optimal [40]. integrity [41]. issues [63].

January [21, 43]. joins [27]. Josephus [47].

keyword [5]. Kirchhoff [3]. knowledge [11]. Kotzig [19].

large [29]. layered [76]. learning [66, 37, 75]. letter [73]. limited [52]. Limiting [1, 39]. load [7]. local [23, 44]. location [31]. locations [31]. lower [77]. lower-bounded [77].

machine [37]. mass [53]. Matching

[69, 62]. Matching-based [69]. matroidal [34]. max [72, 49]. maximum [10]. method [66]. methods [30]. metric [77]. minimization [59]. minimum [25]. minor [61]. mixed [16]. mixed-charging [16]. mobile [18]. mobility [48]. model [68, 44, 17]. Modified [1, 39]. modulo [32]. monadic [46]. monographs [38]. multi [72, 7, 5]. multi-agent [72]. multi-keyword [5]. multi-load [7]. MultiCover [35]. multiplication [32]. mutual [48]. myopic [14]. networks [22, 16, 13, 26, 34, 17, 23, 8]. Newton [66]. Nim [19]. node [17]. notice [39]. number [73]. numerical [30].

oblivious [18, 52]. online [62]. optimal [40, 14]. Optimizing [1, 39]. orientations [13]. orienteering [9]. outliers [79]. overlaps [45].

Packing [28]. parallel [30]. parameterizations [41]. parameters [61]. Parking [18]. partial [2, 30]. partitions [26]. Path [26, 50, 7]. period [8]. phylogenetic [13, 26]. pickup [7]. placements [51]. planar [51]. ply [25]. point [51]. points [25]. policy [72]. polynomials [78]. polyphenyl [3]. Popularity [12]. population [33]. prefix [27]. presence [48]. primal [62]. primal-dual [62]. priority [6]. Privacy [33]. probabilistic [33]. problem [35, 9, 18, 65, 12, 50, 7, 63, 8]. problems [24]. properties [38]. proportional [76]. protocols [33]. pruning [66]. pruning-based [66].

quadruple [63].

Rate [1, 39]. rectangle [76]. red [16]. red-black [16]. reductions [60]. regret [59]. reinforcement [75]. reliability [34, 17]. representations [76]. respect [3]. responses [23]. Restricted [15]. Retraction [39]. robots [18, 52, 14]. Roman [63]. roommate [67, 12]. root [78]. root-finder [78]. rooted [13]. rumor [8]. runs [73].

safety [8]. SARSA [75]. SAT [49]. Scalability [1, 39]. Scale [1, 39]. scenario [2]. scenarios [50]. scheduling [33]. scheme [5]. science [39]. search [64, 5]. security [74]. selection [59]. Self [68]. Self-similarity [68]. sensitive [5]. Services [1, 39]. set [51, 2]. sets [59, 37]. settings [72]. sharing [64, 5]. shortest [50]. side [64]. side-sharing [64]. Sign [74]. Sign-then-encrypt [74]. similarity [68]. sizes [15]. small [29]. social [8]. soft [72]. soft-max [72]. Software [1, 39]. solving [30]. sort [17]. Sorting [27]. space [77]. spanners [16]. spanning [47]. specified [53]. speed [53]. sphere [10]. squares [25, 28]. stability [67]. stimuli [23]. stochastic [8]. strong [59]. Structural [41, 61]. subclasses [4]. subnetwork [17]. Subset [35]. Synergistic [11]. systems [53, 5].

tandems [64]. Theoretical [39]. theory [46]. thief [9]. Time [14]. Tolerance [32]. tolerant [48]. toric [46]. tournaments [29]. transform [73]. transformations [38]. translocation [65]. tree [61]. trees [47, 6]. triangular [14]. Truthful [31]. two [31]. two-facility [31].

unit [25]. used [30]. Using [1, 39].

variants [19]. vehicular [16]. Verifiable [5]. Vertex [47, 41]. Vertex-independent [47]. vertices [51]. via [34]. visibility [52, 48]. voting [40].

way [62]. Wheeler [73]. without [48]. words [45, 46].

References

Latha:2023:OSA

[1] V. L. Padma Latha, N. Sudhakar Reddy, and A. Suresh Babu. Optimizing scalability and availability of cloud based software services using modified scale rate limiting algorithm. Theoretical Computer Science, 943(??):230– 240, January 17, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S0304397522004418. See retraction

notice [39].

Dimant:2025:APS

[2] Shai Michael Dimant and Sven O. Krumke. On approximating partial scenario set cover. Theoretical Computer Science, 1023(??):??, January 1, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S0304397524005085.

Li:2025:EPC

[3] Chengmin Li, Hong Bian, and Haizheng Yu. Extremal polyphenyl chains with respect to the Kirchhoff index. Theoretical Computer Science, 1023(??): ??, January 1, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S0304397524005103.

Panda:2025:ICS

[4] B. S. Panda and Rumki Ghosh. Injective coloring of subclasses of chordal graphs. Theoretical Computer Science, 1023(??):??, January 1, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S0304397524005115.

Zhao:2025:VAB

[5] Jie Zhao, Hejiao Huang, Yongliang Xu, Xiaojun Zhang, Hongwei Du, and Chao Huang. Verifiable attribute-based multikeyword search scheme with sensitive information hiding for cloud-assisted e-healthcare sharing systems. Theoretical Computer Science, 1023(??): ??, January 1, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S0304397524005127.

Yuan:2025:ADT

[6] Jing Yuan and Shaojie Tang. Approximating decision trees with priority hypotheses. Theoretical Computer Science, 1023(??):??, January 1, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S0304397524005139.

Li:2025:DPF

[7] Yifei Li, Ruixi Huang, Hao Ye, Hejiao Huang, and Hongwei Du. Dynamic path finding for multi-load agent pickup and delivery problem. Theoretical Computer Science, 1023(??): ??, January 1, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S0304397524005140.

Zhu:2025:EAS

[8] Jianming Zhu, Ye Xing, Runzhi Li, Smita Ghosh, Priyanshi Garg, and Weili Wu. Efficient algorithm for stochastic rumor blocking problem in social networks during safety accident period. Theoretical Computer Science, 1023(??): ??, January 1, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S0304397524005152.

Bloch-Hansen:2025:ATO

[9] Andrew Bloch-Hansen, Roberto Solis-Oba, and Daniel R. Page. Algorithms for the thief orienteering problem on directed acyclic graphs. Theoretical Computer Science, 1023(??): ??, January 1, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S0304397524005176.

Zhang:2025:AAF

[10] Kaiqi Zhang, Siyuan Zhang, Jirun Gao, Hongzhi Wang, Hong Gao, and Jianzhong Li. Approximation algorithms for finding maximum containing circle and sphere. Theoretical Computer Science, 1023(??): ??, January 1, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S0304397524005188.

Cachin:2025:SK

[11] Christian Cachin, David Lehnherr, and Thomas Studer. Synergistic knowledge. Theoretical Computer Science, 1023(??): ??, January 1, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S030439752400519X.

Ge:2025:PRD

[12] Steven Ge and Toshiya Itoh. Popularity on the roommate diversity problem. Theoretical Computer Science, 1023 (??):??, January 1, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S0304397524005206.

Docker:2025:EFO

[13] Janosch Döcker and Simone Linz. On the existence of funneled orientations for classes of rooted phylogenetic networks. Theoretical Computer Science, 1023(??):??, January 1, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/S0304397524005255.

Goswami:2025:TOG

[14] Pritam Goswami, Avisek Sharma, Satakshi Ghosh, and Buddhadeb Sau. Time optimal gathering of myopic robots on an infinite triangular grid. Theoretical Computer Science, 1023(??): ??, January 1, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/S0304397524005474.

Liu:2025:RHD

[15] Yin Liu, Austen Z. Fan, and Jin-Yi Cai. Restricted Holant dichotomy on domain sizes 3 and 4. Theoretical Computer Science, 1023(??): ??, January 1, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/S0304397524005486.

Bereg:2025:CRB

[16] Sergey Bereg, Yuya Higashikawa, Naoki Katoh, Junichi Teruyama, Yuki Tokuni, and Binhai Zhu. Constructing red-black spanners for mixed-charging vehicular networks. Theoretical Computer Science, 1023(??):??, January 1, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S0304397524005498.

Meng:2025:DDS

[17] Kaiyue Meng and Yuxing Yang. Double declined subnetwork reliability analysis in bubble-sort networks under node fault model. Theoretical Computer Science, 1023(??):??, January 1, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S0304397524005589.

Chakraborty:2025:PPO

[18] Abhinav Chakraborty and Krishnendu Mukhopadhyaya. Parking problem by oblivious mobile robots in infinite grids. Theoretical Computer Science, 1023(??): ??, January 1, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S030439752400570X.

Arun:2025:GKN

[19] Srinivas Arun. Geography, Kotzig’s Nim, and variants. Theoretical Computer Science, 1023(??):??, January 1, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524005747.

Anonymous:2025:EBa

[20] Anonymous. Editorial Board. Theoretical Computer Science, 1023(??): ??, January 1, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S0304397524006091.

Anonymous:2025:Ja

[21] Anonymous. 1 January 2025. Theoretical Computer Science, 1023(??):??, January 1, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Aracena:2025:DED

[22] Julio Aracena, Luis Cabrera-Crot, Adrien Richard, and Lilian Salinas. Dynamically equivalent disjunctive networks. Theoretical Computer Science, 1024(??): ??, January 12, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/S0304397524005164.

Oh:2025:ACR

[23] Shunhao Oh, Dana Randall, and Andréa W. Richa. Adaptive collective responses to local stimuli in anonymous dynamic networks. Theoretical Computer Science, 1024(??): ??, January 12, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/S0304397524005218.

Ibarra:2025:DPC

[24] Oscar H. Ibarra and Ian McQuillan. On decision problems concerning contextual insertions and deletions. Theoretical Computer Science, 1024(??): ??, January 12, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/S030439752400522X.

Durocher:2025:AAM

[25] Stephane Durocher, J. Mark Keil, and Debajyoti Mondal. Approximation algorithms for minimum ply covering of points with unit squares and unit disks. Theoretical Computer Science, 1024(??): ??, January 12, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S0304397524005231.

Lafond:2025:PPP

[26] Manuel Lafond and Vincent Moulton.

Path partitions of phylogenetic networks. Theoretical Computer Science, 1024(??): ??, January 12, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S0304397524005243.

Fertin:2025:SGP

[27] Guillaume Fertin, Géraldine Jean, and Anthony Labarre. Sorting genomes by prefix double-cut-and-joins. Theoretical Computer Science, 1024(??): ??, January 12, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S0304397524005267.

Wu:2025:PSI

[28] Wei Wu, Hiroki Numaguchi, Nir Halman, Yannan Hu, and Mutsunori Yagiura. Packing squares independently. Theoretical Computer Science, 1024(??): ??, January 12, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S0304397524005279.

Yuster:2025:FCS

[29] Raphael Yuster. Finding and counting small tournaments in large tournaments. Theoretical Computer Science, 1024(??): ??, January 12, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/S0304397524005280.

Samaniego:2025:FPD

[30] Cristóbal Samaniego and Guillaume Houzeaux. A formalization of parallel data exchange algorithms used by numerical methods for solving partial differential equations. Theoretical Computer Science, 1024(??):??, January 12, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/S0304397524005292.

Kanellopoulos:2025:TTF

[31] Panagiotis Kanellopoulos, Alexandros A. Voudouris, and Rongsen Zhang. Truthful two-facility location with candidate locations. Theoretical Computer Science, 1024(??):??, January 12, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/S0304397524005309.

Gupta:2025:TAA

[32] Arya Tanmay Gupta and Sandeep S. Kulkarni. Tolerance to asynchrony in algorithms for multiplication and modulo. Theoretical Computer Science, 1024 (??):??, January 12, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S0304397524005310.

Amir:2025:PPP

[33] Talley Amir and James Aspnes. Privacy in population protocols with probabilistic scheduling. Theoretical Computer Science, 1024(??):??, January 12, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524005437.

Li:2025:EHR

[34] Xiao-Yan Li, Zhaoding Lin, Hongbin Zhuang, and Jou-Ming Chang. Enabling high reliability via matroidal connectivity and conditional matroidal connectivity on arrangement graph networks. Theoretical Computer Science, 1024(??): ??, January 12, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S0304397524005449.

Benoist:2025:ESM

[35] Emile Benoist, Guillaume Fertin, and Géraldine Jean. The Exact Subset MultiCover problem. Theoretical Computer Science, 1024(??):??, January 12, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S030439752400553X.

Pin:2025:EDB

[36] Jean-Éric Pin. Euclidean division by d in base b. Theoretical Computer Science, 1024(??):??, January 12, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/S0304397524005541.

Singh:2025:CML

[37] Barkha Singh, Sreedevi Indu, and Sudipta Majumdar. Comparison of machine learning algorithms for classification of Big Data sets. Theoretical Computer Science, 1024(??): ??, January 12, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/S0304397524005553.

delaTour:2025:APT

[38] Thierry Boy de la Tour. Algebraic properties and transformations of monographs. Theoretical Computer Science, 1024(??): ??, January 12, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/S0304397524005565.

Latha:2025:RNO

[39] V. L. Padma Latha, N. Sudhakar Reddy, and A. Suresh Babu. Retraction notice to “Optimizing Scalability and Availability of Cloud Based Software Services Using Modified Scale Rate Limiting Algorithm” [Theoretical computer science 943 (2023) 230–240]. Theoretical Computer Science, 1024(??):??, January 12, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/S0304397524005619. See [1].

Chierichetti:2025:IOI

[40] Flavio Chierichetti. Instance-optimal information-based voting. Theoretical Computer Science, 1024(??):??, January 12, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S0304397524005620.

Gima:2025:SPV

[41] Tatsuya Gima, Tesshu Hanaka, Yasuaki Kobayashi, Ryota Murai, Hirotaka Ono, and Yota Otachi. Structural parameterizations of vertex integrity. Theoretical Computer Science, 1024(??): ??, January 12, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S0304397524005711.

Anonymous:2025:EBb

[42] Anonymous. Editorial Board. Theoretical Computer Science, 1024(??): ??, January 12, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S0304397524006170.

Anonymous:2025:Jb

[43] Anonymous. 12 January 2025. Theoretical Computer Science, 1024(??):??, January 12, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Delporte-Gallet:2025:DCA

[44] Carole Delporte-Gallet, Hugues Fauconnier, Pierre Fraigniaud, and Mikaël Rabie. Distributed computing in the asynchronous LOCAL model. Theoretical Computer Science, 1025(??): ??, February 2, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S0304397524005693.

Anselmo:2025:DAW

[45] Marcella Anselmo, Manuela Flores, and Maria Madonia. Density of k-ary words with 0, 1, 2-error overlaps. Theoretical Computer Science, 1025(??): ??, February 2, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S0304397524005759.

Berthe:2025:MTT

[46] Valérie Berthé, Toghrul Karimov, Joris Nieuwveld, Joël Ouaknine, Mihir Vahanwala, and James Worrell. The monadic theory of toric words. Theoretical Computer Science, 1025(??): ??, February 2, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S0304397524005760.

He:2025:VIS

[47] Qi He, Yan Wang, Jianxi Fan, and Baolei Cheng. Vertex-independent spanning trees in complete Josephus cubes. Theoretical Computer Science, 1025(??): ??, February 2, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S0304397524005863.

Pramanick:2025:FTM

[48] Subhajit Pramanick, Saswata Jana, and Partha Sarathi Mandal. Fault-tolerant mutual visibility without any axis agreement in presence of mobility failure. Theoretical Computer Science, 1025(??):

??, February 2, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S0304397524005875.

Panolan:2025:MSC

[49] Fahad Panolan and Hannane Yaghoubizade.

On MAX–SAT with cardinality constraint. Theoretical Computer Science, 1025(??):??, February 2, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S0304397524005887.

Kasperski:2025:ASP

[50] Adam Kasperski and Paweł Zieliński. Approximating the shortest path problem with scenarios. Theoretical Computer Science, 1025(??):??, February 2, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524005899.

Damaschke:2025:CPN

[51] Peter Damaschke, Fredrik Ekstedt, and Raad Salman. On central placements of new vertices in a planar point set. Theoretical Computer Science, 1025(??): ??, February 2, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S0304397524005905.

DiLuna:2025:GCL

[52] Giuseppe Antonio Di Luna, Ryuhei Uehara, Giovanni Viglietta, and Yukiko Yamauchi. Gathering on a circle with limited visibility by anonymous oblivious robots. Theoretical Computer Science, 1025(??):??, February 2, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/S0304397524005917.

Anderson:2025:CMA

[53] David F. Anderson and Badal Joshi. Chemical mass–action systems as analog computers: Implementing arithmetic computations at specified speed. Theoretical Computer Science, 1025(??): ??, February 2, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/S0304397524006005.

Bensmail:2025:ADC

[54] Julien Bensmail, Hervé Hocquard, and Clara Marcille. Adding direction constraints to the 1–2–3 Conjecture. Theoretical Computer Science, 1025(??): ??, February 2, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/S0304397524006029.

Anonymous:2025:EBc

[55] Anonymous. Editorial Board. Theoretical Computer Science, 1025(??): ??, February 2, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/S0304397524006340.

Anonymous:2025:Fa

[56] Anonymous. 2 February 2025. Theoretical Computer Science, 1025(??):??, February 2, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2025:Fb

[57] Anonymous. 12 February 2025. Theoretical Computer Science, 1026(??):??, February 12, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2025:EBd

[58] Anonymous. Editorial Board. Theoretical Computer Science, 1026(??): ??, February 12, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https://www.

sciencedirect.com/science/article/
pii/S0304397524006595.

Guo:2025:TSR

[59] Hongjie Guo, Jianzhong Li, and Hong Gao. Towards strong regret minimization sets: Balancing freshness and diversity in data selection. Theoretical Computer Science, 1026(??):??, February 12, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https://www.

sciencedirect.com/science/article/
pii/S0304397524006030.

Abboud:2025:FGR

[60] Elie Abboud and Noga Ron-Zewi.

Finer-grained reductions in fine-grained hardness of approximation. Theoretical Computer Science, 1026(??):??, February 12, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https://www.

sciencedirect.com/science/article/
pii/S0304397524005930.

Gima:2025:DTM

[61] Tatsuya Gima, Soh Kumabe, Kazuhiro Kurita, Yuto Okada, and Yota Otachi.

Dichotomies for tree minor containment with structural parameters. Theoretical Computer Science, 1026(??): ??, February 12, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https://www.

sciencedirect.com/science/article/
pii/S0304397524006017.

Kakimura:2025:DPD

[62] Naonori Kakimura and Tomohiro Nakayoshi. Deterministic primal-dual algorithms for online k-way matching with delays. Theoretical Computer Science, 1026(??): ??, February 12, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https://www.

sciencedirect.com/science/article/
pii/S0304397524006054.

Palagiri:2025:CIC

[63] Venkata Subba Reddy Palagiri, Guru Pratap Sharma, and Ismael G. Yero. Complexity issues concerning the quadruple Roman domination problem in graphs. Theoretical Computer Science, 1026(??): ??, February 12, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https://www.

sciencedirect.com/science/article/
pii/S0304397524006303.

Marcus:2025:EIS

[64] Shoshana Marcus, Dina Sokol, and Sarah Zelikovitz. Exact and inexact search for 2d side-sharing tandems. Theoretical Computer Science, 1026(??): ??, February 12, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https://www.

sciencedirect.com/science/article/
pii/S0304397524006224.

Constantin:2025:EAA

[65] Maria Constantin and Alexandru Popa.

Exact and approximation algorithms for the contiguous translocation distance problem. Theoretical Computer Science, 1026(??):??, February 12, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https://www.

sciencedirect.com/science/article/
pii/S0304397524006200.

Chen:2025:APB

[66] Shuzhen Chen, Yuan Yuan, Youming Tao, Tianzhu Wang, Zhipeng Cai, and Dongxiao Yu. Adaptive pruning-based Newton’s method for distributed learning. Theoretical Computer Science, 1026 (??):??, February 12, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https://www.

sciencedirect.com/science/article/
pii/S0304397524006042.

Bonifaci:2025:ERA

[67] Vincenzo Bonifaci and Helena Rivera Dallorto. Egalitarian roommate allocations: Complexity and stability. Theoretical Computer Science, 1026(??): ??, February 12, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https://www.

sciencedirect.com/science/article/
pii/S0304397524006261.

Barrett:2025:SSC

[68] Jordan Barrett, Bogumił Kamiński, Paweł Prałat, and François Théberge. Self-similarity of communities of the ABCD model. Theoretical Computer Science, 1026(??):??, February 12, 2025. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397524006297.

Kobayashi:2025:EAI

[69] Yusuke Kobayashi, Ryoga Mahara, and Souta Sakamoto. EFX allocations for indivisible chores: Matching-based approach. Theoretical Computer Science, 1026(??):??, February 12, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https://www.

                                        sciencedirect.com/science/article/
                                        pii/S0304397524006273.

Anonymous:2025:Fc

[70] Anonymous. 19 February 2025. Theoretical Computer Science, 1027(??):??, February 19, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2025:EBe

[71] Anonymous. Editorial Board. Theoretical Computer Science, 1027(??): ??, February 19, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https://www.

                                        sciencedirect.com/science/article/
                                        pii/S0304397524006765.

Bernasconi:2025:EDS

[72] Martino Bernasconi, Federico Cacciamani, Simone Fioravanti, Nicola Gatti, and Francesco Trovò. The evolutionary dynamics of soft-max policy gradient in multi-agent settings. Theoretical Computer Science, 1027(??): ??, February 19, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https://www.

                                        sciencedirect.com/science/article/
                                        pii/S0304397524006285.

Biagi:2025:NEL

[73] Elena Biagi, Davide Cenzato, Zsuzsanna Lipták, and Giuseppe Romana. On the number of equal-letter runs of the bijective Burrows–Wheeler transform. Theoretical Computer Science, 1027(??): ??, February 19, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https://www.

sciencedirect.com/science/article/
pii/S0304397524006212.

Chen:2025:STE

[74] Yanbo Chen and Yunlei Zhao. Signthen-encrypt with security enhancement and compressed ciphertext. Theoretical Computer Science, 1027(??): ??, February 19, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https://www.

sciencedirect.com/science/article/
pii/S0304397524006236.

Yao:2025:ISD

[75] Guangyu Yao, Nan Zhang, Zhenhua Duan, and Cong Tian. Improved SARSA and DQN algorithms for reinforcement learning. Theoretical Computer Science, 1027(??):??, February 19, 2025. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https://www.

sciencedirect.com/science/article/
pii/S030439752400642X.

Haase:2025:LAP

[76] Carolina Haase and Philipp Kindermann. On layered area-proportional rectangle contact representations. Theoretical Computer Science, 1027(??): ??, February 19, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https://www.

                                        sciencedirect.com/science/article/
                                        pii/S0304397524006388.

Liang:2025:DAL

[77] Ting Liang, Xiaoliang Wu, Jinhui Xu, and Qilong Feng. The distributed algorithms for the lower-bounded kcenter clustering in metric space. Theoretical Computer Science, 1027(??): ??, February 19, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https://www.

                                        sciencedirect.com/science/article/
                                        pii/S0304397524005929.

Pan:2025:NFR

[78] Victor Y. Pan, Soo Go, Qi Luan, and Liang Zhao. A new fast rootfinder for black box polynomials. Theoretical Computer Science, 1027(??): ??, February 19, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https://www.

                                        sciencedirect.com/science/article/
                                        pii/S030439752400639X.

Dabas:2025:FAC

[79] Rajni Dabas, Neelima Gupta, and Tanmay Inamdar. FPT approximation for capacitated clustering with outliers. Theoretical Computer Science, 1027(??): ??, February 19, 2025. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https://www.

                                        sciencedirect.com/science/article/
                                        pii/S0304397524006431.